# Anton Alexa

### v4

@@ Conclusion
We have proved a quantitative boundary delocalization principle for Dirichlet eigenfunctions on smooth strictly convex domains, showing that no individual high–frequency eigenmode can asymptotically dominate the boundary energy of a sublinear spectral packet. The result excludes persistent boundary concentration at the level of single eigenfunctions when compared to neighboring modes in vanishing relative frequency windows.
The main mode–to–packet estimate is entirely static and relies only on the Rellich identity. For the multi–mode bias exclusion we additionally use the boundary local Weyl law; integrated boundary Weyl asymptotics for the Dirichlet Laplacian [2, 4, 8] are recorded for context. The argument is robust under eigenvalue crossings and depends only on global spectral information. This structural invariance makes the packet framework natural for deformation-type questions, where multiplicities and mode labeling may vary.
From a conceptual viewpoint, the result provides a mode–to–packet comparison mechanism on the boundary, analogous to bulk delocalization phenomena but valid in a purely deterministic and geometric setting. For single–mode comparison, the sublinear packet regime with $N_{k}\to\infty$ and $N_{k}=o(k)$ is sufficient. Notably, the single–mode packet delocalization estimate is obtained without using local Weyl asymptotics. For packet zero–mean cancellation and multi–mode bias exclusion, the present method identifies the threshold scale $N_{k}\gg k^{1-2/d}$. The packet–level perspective developed here is particularly suited to averaged spectral functionals, where single–eigenfunction control is often too weak. In addition, we obtain a multi–mode delocalization principle: no positive fraction of modes in a sublinear packet can maintain strong boundary correlation with a fixed zero–mean weight.
Several directions for further investigation suggest themselves. First, it would be natural to extend the argument to other boundary conditions or to more general elliptic operators with boundary. Second, one may ask whether analogous packet–level delocalization phenomena hold on nonconvex domains or on Riemannian manifolds with boundary, where boundary Weyl remainders are less rigid. Finally, the present comparison principle may serve as a useful input in the study of boundary spectral statistics and correlations, complementing existing microlocal and dynamical approaches.

### v5

@@ Conclusion
The principal new contribution of this paper is the identification of a threshold hierarchy for boundary correlation cancellation of Dirichlet eigenfunctions on smooth strictly convex domains. The minimal packet scale required for the energy–weighted average of boundary correlation coefficients to vanish is determined by the vanishing order of curvature moments of the weight with respect to domain geometry. Theorems 5.7 and 5.10 establish two levels of this hierarchy, via the boundary local Weyl law [2, 4, 8]: (i) If $\int_{\partial\Omega}w\,d\sigma=0$, packet zero–mean cancellation holds at threshold $N_{k}\gg k^{1-2/d}$. (ii) If additionally $\int_{\partial\Omega}H\,w\,d\sigma=0$, the threshold weakens to $N_{k}\gg k^{(d-3)/d}$, which for $d=3$ reduces to $N_{k}\to\infty$. This hierarchy, stratified by the vanishing order of boundary curvature moments of $w$, is a new structural feature of the boundary spectral problem with no analogue in bulk delocalization results.
As a structural consequence of the Rellich identity alone, Theorem 5.1 establishes that the ratio of the boundary energy of a single eigenmode to the packet sum satisfies $c/N_{k}\leq E_{k}/\sum_{m=k}^{k+N_{k}-1}E_{m}\leq C/N_{k}$. The rate $1/N_{k}$ is sharp and follows from global Rellich bounds alone, without any use of Weyl asymptotics or microlocal analysis.
The argument is robust under eigenvalue crossings and depends only on global spectral information. This structural invariance makes the packet framework particularly natural for deformation–type questions, where multiplicities and mode labeling may vary.
Several directions for further investigation suggest themselves. First, it would be natural to extend the argument to other boundary conditions or to more general elliptic operators with boundary. Second, one may ask whether analogous packet–level delocalization phenomena hold on nonconvex domains or on Riemannian manifolds with boundary, where boundary Weyl remainders are less rigid. Third, the sharpness of the enhanced threshold $k^{(d-3)/d}$ under the double vanishing moment condition, and of the threshold hierarchy in higher dimensions, remains to be investigated. Finally, the present comparison principle may serve as a useful input in the study of boundary spectral statistics and correlations, complementing existing microlocal and dynamical approaches.